### v1

# Copper delocalization leads to ultralow thermal conductivity in chalcohalide CuBiSeCl2

Abstract
Mixed anion halide-chalcogenide materials have attracted considerable attention due to their exceptional optoelectronic properties, making them promising candidates for various applications. Among these, CuBiSeCl2 has recently been experimentally identified with remarkably low lattice thermal conductivity ($\kappa_{L}$). In this study, we employ Wigner transport theory combined with neuroevolution machine learning potential (NEP)-assisted self-consistent phonon calculations to unravel the microscopic origins of this low $\kappa_{L}$. Our findings reveal that the delocalization and weak bonding of copper atoms are key contributors to the strong phonon anharmonicity and wavelike tunneling (random walk diffusons). These insights deepen our understanding of the relationship between bonding characteristics, anharmonicity, delocalization, and vibrational dynamics, paving the way for the design and optimization of CuBiSeCl2 and analogous materials for advanced phonon engineering applications.
keywords:
Anharmonic lattice dynamics; Delocalization of copper; Lattice thermal conductivity; Mixed-anion halide-chalcogenide
Introduction
Mixed-anion compounds, such as halide-chalcogenides, have recently gained attention as promising candidates for thermoelectric and photovoltaic applications due to their ability to achieve low thermal conductivity while maintaining favorable electronic properties 1, 2. These compounds exhibit increased structural anisotropy, dimensional changes, and the formation of novel structural types 3. The structural modifications induced by variations in bonding promote the emergence of new physical phenomena and offer opportunities to tune both optoelectronic and thermoelectronic properties by altering the electronic structure 4, 5, 6, 7, 8.
CuBiSeCl2, a mixed-metal chalcohalide, stands out for its potential to exhibit both low thermal conductivity and favorable electronic properties. Previous experimental studies have shown that CuBiSeCl2 has a band gap of 1.33 eV, as measured by diffuse reflectance UV-vis spectroscopy, making it suitable for photovoltaic applications 3. Additionally, it demonstrates a low thermal conductivity of 0.27 W m-1 K-1 at room temperature 3. However, the complex crystal structure has limited comprehensive studies on the underlying mechanisms responsible for its low $\kappa_{L}$ , particularly for such intricate mixed-anion compounds.
In this study, we investigate the microscopic mechanisms underlying the low $\kappa_{L}$ of CuBiSeCl2 by combining machine learning potentials with anharmonic lattice dynamics. We integrate machine learning potentials with the Boltzmann transport equation (BTE) and molecular dynamics (MD) simulations to conduct a detailed analysis of phonon transport and scattering mechanisms. This computational framework allows for more accurate predictions of thermal conductivity and provides deeper insights into the phonon anharmonicity and glassy thermal transport in crystalline compounds. Our results not only elucidate the origins of the low $\kappa_{L}$ in CuBiSeCl2 but also offer a new strategy for further reducing $\kappa_{L}$ through the delocalization of atoms and the introduction of softer chemical and vibrational bonds. The computational details can be found in the Supplemental Material.
Results and discussion
A. Crystal structure
Figure 1: (a) The conventional cell structure of CuBiSeCl2. The yellow, red, blue and purple atoms represent Cu, Bi, Se, and Cl, respectively. (b) The temperature-dependent phonon spectrum from T = 0 K to 700 K for CuBiSeCl2, along the high-symmetry points which are defined as $\Gamma$ (0, 0, 0), X (0.5, 0, 0), S (0.5, 0.5, 0), Y (0, 0.5, 0), Z (0, 0, 0.5), U (0.5, 0, 0.5), R (0.5, 0.5, 0.5) and T (0, 0.5, 0.5). (c) Projected phonon density of states (PDOS) of CuBiSeCl2 at 0 K, 300 K, and 700 K. The atomic participation ratio (APR) of Cu (d) and Bi (e) atoms projected onto the phonon bands at T = 0 K. The color bar represents the atomic participation ratio (APR) for the specified atom in the phonon modes. An APR of 1 (red) indicates that the atom is highly involved in the corresponding vibrational mode, whereas an APR of 0 (blue) signifies little participation of the atom in that mode. (f) Calculated spectral and cumulative particle-like lattice thermal conductivity ($\kappa_{p}$) of CuBiSeCl2 using the SCPH+3,4ph model at 300 K in different axes.
The crystal structure of CuBiSeCl2 belongs to space group $Pnma$ with number of 62 and the optimized lattice constants, using PBE potentials, are $a$ = 8.73 Å, $b$ = 4.05 Å, and $c$ = 12.99 Å, in which the van der Waals (vdW) interactions are described by the optB86b functional 9, 10. These values are consistent with the experimental values, which are $a$ = 8.78 Å, $b$ = 4.00 Å, and $c$ = 13.14 Å 3. As illustrated in Fig. 1(b), the phonon spectrum of the PBE functional is free of imaginary frequencies, indicating the dynamical stability of CuBiSeCl2. The primitive cell comprises 20 atoms: four Bi, four Se, four Cu, and eight Cl respectively.
B. Anharmonic lattice dynamics
The impact of fourth-order anharmonicity on the phonon spectra at finite temperatures is depicted in Fig. 1(b), using the interatomic force constants (IFCs) obtained from MLP. The overall phonon hardening is observed in crystalline CuBiSeCl2 as the temperature increases from 0 to 700 K. This suggests the presence of strong lattice anharmonicity in CuBiSeCl2 and highlights the importance of properly treating temperature renormalization 11, 12, 13, 14, 15. In particular, the low-frequency phonons below 3 Thz are predominantly influenced by the Cu (yellow) and Bi (red) atoms, as evidenced by the projected phonon density of states (PDOS) in Fig. 1(c).
Furthermore, atom-decomposed partial DOS in Fig. 1(c) and atomic participation ratios in Figs. 1(d-e) show that Cu and Bi atoms contribute mostly to the low-lying flattened phonon bands in the frequency lower than 2.0 THz. This suggests that strong phonon scattering is occurring, which suppresses the primary heat carriers, i.e., acoustic phonon modes 16, 17, 18. In contrast, the high-frequency phonon modes around 6 THz in CuBiSeCl2 are predominantly contributed by Se and Cl atoms and exhibit a temperature-dependent softening in Fig. 1(c). Cumulated $\kappa_{p}$ shown in Fig. 1(f) also approve that low frequency phonons less than 1.5 THz are the main contributor to the $\kappa_{p}$ of CuBiSeCl2.
C. Delocalization of copper atoms
Fig. 2(a) shows the mean square displacements (MSDs) of atoms in CuBiSeCl2 as a function of temperature along the $b$-axis. MSD is a measure of the deviation of the position of an atom with respect to a reference position over time. It indicates the atomic displacements of Cu are significantly larger than those of other atoms Bi, Se, and Cl, suggesting that the bonding of Cu atoms is heavily delocalized and soft 11.
In Fig. 2(b), the displacement of Cu atoms shows a relatively deep energy well with a flat bottom as a function of the phonon mode along the $b$-axis. The U-shaped potential energy curve significantly deviates from the harmonic approximation (quadratic fit) shown as the dotted line in Fig. 2(b). The U-shaped potential energy of Cu atom displacement can be well fitted by a fourth-order polynomial (solid yellow line in Fig. 2(b)), expressed as $y=1.41x^{4}+0.38x^{2}$, indicating strong quartic phonon anharmonicity.
The red color corresponds to complete localization and blue corresponds to complete delocalization of electrons in 2D electron localization function (ELF) of CuBiSeCl2 shown in Fig. 2(c). It can be seen that the localized electronic density around Cu is extremely low, significantly lower than the other atoms, which reveals that the Se and Cl elements exhibit a much stronger attraction of electrons than the Cu elements, suggesting a greater degree of delocalization for Cu atoms.
The $\kappa_{L}$ of CuBiSeCl2 at 300 K in Fig. 2(d-f), calculated using homogeneous nonequilibrium molecular dynamics (HNEMD) is the averaged result from five 2 ns simulation runs, totally 10 ns 19, 20, 21. We employed precise NEP to obtain the intrinsic $\kappa_{L}$ of CuBiSeCl2, highlighted in red color in Fig.2(d-f). The thicker solid line corresponds to the averaged result, and the converged $\kappa_{L}$ values are marked on the right side of the figure. These values indicate a significantly lower intrinsic $\kappa_{L}$ along the $c$-axis (0.34 W m-1 K-1) compared to the $a$- and $b$-axes (0.50 and 0.55 W m-1 K-1), consistent with the final result shown in Fig. 4(d) 13, 22.
Based on Fig. 2(a-c), we observe a significant difference in the delocalization of Cu compared to Bi, Se, and Cl. To investigate the impact of the delocalization of different atoms on $\kappa_{L}$, we used the same HNEMD method to calculate $\kappa_{L}$ along $a$-, $b$- and $c$-axes of the system after fixing different types of atoms. The results are shown in Fig. 2(d-f). Yellow, purple, and blue represent the fixed atoms of Cu, Bi, and Se, respectively. In all three directions, for the cases where Bi and Se atoms are fixed, a significant decrease in $\kappa_{L}$ compared to the intrinsic values is observed, indicating that the strong localization of Bi and Se atoms enhances the phonon transport of CuBiSeCl2. However, fixing Cu in the $a$- and $b$-axes reveals its counterintuitive phenomenon, as its fixation leads to an increase in $\kappa_{L}$ compared to the intrinsic value by 0.08 and 0.67 W m-1 K-1, respectively. This highlights that the delocalization of Cu has a negative impact on $\kappa_{L}$, especially on the $b$-axis 23.
Figure 2: (a) The mean square atomic displacements (MSDs) with temperatures for CuBiSeCl2. (b) Potential energy per atom with displacement of Cu (yellow) atoms along the $b$-axis direction with quartic and quadratic polynomial fitting. It writes as $y=1.41x^{4}+0.38x^{2}$ and $y=1.58x^{2}$. For the Bi (red), Se (blue), and Cl (purple) atoms, the quartic polynomial fittings along the $b$-axis direction are as follows: $y=0.67x^{4}+1.75x^{2}$, $y=1.20x^{4}+1.39x^{2}$ and $y=0.015x^{4}+2.76x^{2}$. The cubic term is minimal and less than $10^{-5}$ 24, 22. (c) Electron localization function (ELF) of CuBiSeCl2. The red color corresponds to complete localization and the blue color corresponds to complete delocalization of electrons. $\kappa_{L}$ predicted by the HNEMD method with NEP vs integration time for CuBiSeCl2 in $a$-axis (d), $b$-axis (e), and $c$-axis (f) respectively. The values labeled on the right represent the averaged final converged $\kappa_{L}$.
D. Thermal transport properties
Based on Peierls-Boltzmann transport theory, the lattice thermal conductivity $\kappa_{p}$ is given by,
$\kappa_{p}=\frac{\hbar^{2}}{k_{B}T^{2}VN_{0}}\sum_{\lambda}n_{\lambda}(n_{\lambda}+1)\omega^{2}_{\lambda}\bm{v}\lambda\otimes\bm{v}\lambda\bm{\tau}_{\lambda},$ (1)
and $\kappa_{c}$ describing tunneling and wavelike of coherent phonons, governed by the Wigner transport equation, is defined as 25, 13,
$\kappa_{c}=$ $\frac{\hbar^{2}}{k_{B}T^{2}VN_{0}}\sum_{\bm{q}}\sum_{s\neq s^{\prime}}\frac{\omega_{\bm{q}}^{s}+\omega_{\bm{q}}^{s^{\prime}}}{2}\bm{v}_{\bm{q}}^{s,\bm{s}^{\prime}}\bm{v}_{\bm{q}}^{\bm{s}^{\prime},s}$ (2)
×$\times\frac{\omega_{\bm{q}}^{s}n_{\bm{q}}^{s}\left(n_{\bm{q}}^{s}+1\right)+\omega_{\bm{q}}^{s^{\prime}}n_{\bm{q}}^{s^{\prime}}\left(n_{\bm{q}}^{s^{\prime}}+1\right)}{4\left(\omega_{\bm{q}}^{s^{\prime}}-\omega_{\bm{q}}^{s}\right)^{2}+\left(\Gamma_{\bm{q}}^{s}+\Gamma_{\bm{q}}^{s^{\prime}}\right)^{2}}$
×$\times\left(\Gamma_{\bm{q}}^{s}+\Gamma_{\bm{q}}^{s^{\prime}}\right),$
The total $\kappa_{L}$ = $\kappa_{p}$ + $\kappa_{c}$, and the parameters in these equations can be found in the Supplementary Material.
From Eq. (1) and Eq. (2), $\kappa_{L}$ is a result of coupling between the harmonic and the anharmonic terms. To further elucidate the microscopic origins of ultralow $\kappa_{L}$, we analyzed several critical parameters influencing $\kappa_{L}$, such as heat capacity, phonon velocities, scattering rates, Grüneisen parameter, weighted scattering phase space, and lifetimes of CuBiSeCl2, as shown in Fig. 3.
Heat capacity ($C_{V}$) is $3Nk_{B}$ in a typical solid at high temperature (the Dulong–Petit limit), where $N$ is the number of particles and $k_{B}$ is the Boltzmann constant. Here in Fig. 3(a), with temperature going higher than 300 K, the $C_{V}$ given out by DFT calculation is very close to the theoretical value $3Nk_{B}$ = 0.295 J g-1 K-1. Also, this value of CuBiSeCl2 is even lower than 0.36 J g-1 K-1 of liquid-like thermoelectric Cu2Se 26 and a little higher than 0.156 J g-1 K-1 of renowned PbTe 27.
[…]
The extent of anharmonicity is typically measured by the Grüneisen parameter, illustrated in Fig. 3(c). Below 1.5 THz, there is an intertwinement between the acoustic and optical phonon branches at 300 K. This is closely linked to the Cu and Bi atoms and the bonding hierarchy. Moreover, the total Grüneisen value is 1.50, close to PbTe value of 1.45 31, indicating CuBiSeCl2 exhibits substantial anharmonicity.
To deepen the analysis of phonon scattering processes, we computed the phonon scattering phase space, which evaluates the channel available for phonon-phonon scattering. As shown in Fig. 3(d), the weighted phase space for both three-phonon (3ph) and four-phonon (4ph) processes rises with temperature, from 300 K to 700 K. Notably, the 4ph phase space exceeds that of 3ph processes, leading to a significant contribution of 4ph scattering in CuBiSeCl2. The peak in the 4ph phase space around 1.5 THz aligns with the partial density of states (PDOS) for Cu atoms, shown in Fig. 1(c). The presence of low-energy flattening modes enhances 4ph scattering channels, which elevates the 4ph scattering rates and, consequently, further limits thermal transport in CuBiSeCl2.
We evaluate the 3ph scattering channels involving three acoustic phonons (AAA, purple), two acoustic phonons combining with one optical phonon (AAO, red), and one acoustic phonon combining with two optical phonons (AOO, orange), as shown in Fig. 3(e). The scattering process of the AAO mode is found to be as strong as the AOO mode below 1 THz, meaning that the AAO and AOO phonons play dominant roles in heat transport. Also, as temperature increases to 700 K in Fig. 3(e) the scattering process of AAA mode becomes much stronger compared to other modes, the maximum scattering rates are changed from around $10^{-4}$ to $10^{-3}$ ps-1, which significantly reduce the $\kappa_{L}$ of CuBiSeCl2 at higher temperatures.
Figure 3: Microscopy heat transport parameters of CuBiSeCl2. (a) Specific heat capacity at constant volume ($C_{V}$). (b) Squared phonon group velocities ($v^{2}$) in the harmonic approximation and anharmonic renormalization approximation at 300 K. (c) Grüneisen parameter. (d) Weighted phonon scattering phase space of 3ph and 4ph at 300 K and 700 K. (e) Phonon scattering rates of 3ph at 300 K and 700 K. (f) Phonon lifetime of at 300 K. The red horizontal line represents the Wigner limit, expressed as $\tau_{\text{threshold}}=[\Delta\omega_{\text{avg}}]^{-1}$, where $\Delta\omega_{\text{avg}}$ is the average phonon inter-band spacing 32. $\Delta\omega_{\text{avg}}$ is determined by $\omega_{\text{max}}/3N_{\text{at}}$, with $N_{\text{at}}$ being the number of atoms per unit cell and $\omega_{\text{max}}$ being the highest frequency. Phonons with lifetimes longer than $\tau_{\text{threshold}}$ primarily affect thermal conductivity via population mechanisms, while shorter lifetimes contribute to wavelike tunneling. The black curve shows the Ioffe-Regel limit, $\tau_{\text{IR}}=2\pi/\omega$, in which $\omega$ is the phonon frequency. Phonons lifetime exceeding $\tau_{\text{IR}}$ can be described by the Wigner transport equation 33, 15, 22.
Recently, a seminal concept is introduced known as the Wigner limit in time, which has been applied to distinguish phonons into various thermal transport regimes, primarily consisting of particle-like propagation and wavelike tunneling 32, 25, 34. The Wigner limit in time is characterized as $\tau_{\text{threshold}}=3N_{\text{at}}/\omega_{\text{max}}$, where $N_{\text{at}}$ represents the number of atoms in the unit cell, and $\omega_{\text{max}}$ denotes the maximum phonon frequency. According to this criterion, phonons with lifetimes exceeding $\tau_{\text{threshold}}$ exhibit particle-like behavior and are the main contributors to the $\kappa_{p}$ from phonon population. In contrast, those with shorter lifetime exhibit wavelike tunneling behavior, primarily contributing to the $\kappa_{c}$ from coherence phonons.
For the case of CuBiSeCl2, after accounting for both 3ph and 4ph scattering mechanisms, most phonons have lifetimes below the Wigner limit at 300 K, underscoring the significance of wavelike tunneling as a transport channel, as illustrated in Fig. 3(f). This suggests that the coherence contribution from wavelike tunneling plays an important role in the overall thermal conductivity of CuBiSeCl2, as observed in Fig. 4.
Notably, when 4ph scattering is incorporated into SCPH+3,4ph model, the particle-like propagation channel is notably suppressed compared with SCPH+3ph model, making the wavelike tunneling channel more important in thermal transport. The findings illustrate the limitations of traditional phonon-gas models in explaining the thermal behavior of CuBiSeCl2 and emphasize the need for advanced methods, such as the SCPH+3,4ph together with $\kappa_{c}$, to more accurately capture the underlying heat transfer mechanisms in this material.
We further calculated the temperature-dependent $\kappa_{L}$ across four theoretical models to examine the effects of phonon energy renormalization and multi-phonon interactions arising from cubic and quartic phonon anharmonicities, as shown in Fig. 4(a). Results reveal that $\kappa_{L}$ exhibits anisotropy in Fig. 4(b), with the $\kappa_{p}$ along the $a$- and $b$-axes being approximately 100% and 160% higher, respectively, than along the $c$-axis. The lowest level of thermal transport theory, HA+3ph model gives $\kappa_{p,HA}^{3ph}$ values in the $c$-axis direction of 3.08, 0.34, and 0.14 W m-1 K-1 at 50, 300, and 700 K, respectively.
As previously noted, anharmonic phonon renormalization plays a critical role in predicting finite-temperature $\kappa_{L}$ in complex compounds. By including phonon energy shifts in Fig. 4(b), we extend the HA+3ph model to the more precise SCPH+3ph model, yielding $\kappa_{p,SCPH}^{3ph}$ values of 3.67, 0.56, and 0.33 W m-1 K-1 in the $c$-axis at 50, 300, and 700 K, respectively 19%, 67%, and 129% higher than $\kappa_{p,HA}^{3ph}$, indicating significant anharmonic phonon renormalization in CuBiSeCl2 that intensifies with temperature. This trend aligns with Fig. 1(b), which shows pronounced phonon frequency shifts as temperature increases.
Moreover, quartic anharmonicity not only induces substantial phonon frequency shifts but also results in pronounced 4ph scattering. We obtained the average $\kappa_{p,SCPH}^{3,4ph}$ values of 4.24, 0.58, and 0.32 W m-1 K-1 at 50, 300, and 700 K, respectively, shown in Fig. 4(b). These values represent reductions of 17%, 33%, and 42% compared to $\kappa_{p,SCPH}^{3ph}$, highlighting the impact of 4ph scattering 13, 18, 35. This result confirms that phonon hardening and 4ph scattering, driven by quartic anharmonicity, have counteractive effects on $\kappa_{L}$.
Figure 4: (a) A schematic illustration demonstrating the prediction accuracy of $\kappa_{L}$ across different hierarchical thermal transport methods, including the HA+3ph, SCPH+3ph, SCPH+3,4ph, and SCPH+3,4ph+$\kappa_{c}$, where 3ph indicates only three-phonon scattering is included and 3,4ph means both three-phonon and four-phonon scattering are considered. HA means harmonic approximation, and SCPH is short for self-consistent phonon. (b) $\kappa_{p}$ of CuBiSeCl2 using HA+3ph, SCPH+3ph, SCPH+3,4ph approaches. (c) $\kappa_{p}$ of CuBiSeCl2 compare with $\kappa_{c}$. (d) $\kappa_{L}$ of CuBiSeCl2 using BTE method and MD method compare with experimental values 3.
Nevertheless, the coherence contributions are not negligible as shown in Fig. 4(c). By further considering the off-diagonal terms of heat flux operators, the total $\kappa_{L}$ significantly enhanced by incorporating both $\kappa_{p,SCPH}^{3,4ph}$ and $\kappa_{c}$ 25. The ratio $\kappa_{c}/\kappa_{p}$ (red star) is greater than 0.5 at 500 K and exhibits a pronounced increase with temperature. Specifically, in the $a$-axis direction, the ratio of $\kappa_{c}/\kappa_{p}$ (blue cubic) reaches 0.94 at 700 K, which is 1.7 times larger than the 0.55 observed at 300 K. This observation suggests that the contribution of $\kappa_{c}$ is equally important as $\kappa_{p}$ in crystalline CuBiSeCl2 especially in $a$-axis and high temperature. The dominant role of coherence contribution in thermal transport was also observed in other complex compounds such as Cs2AgBiBr6 33, Cu12Sb4S13 36, CsPbBr3 37, TlAgI2 15, La2Zr2O7 32, Tl9SbTe6 22, and even simple cubic CsCl 35.
We further use the HNEMD method to assess the intrinsic $\kappa_{L}$ of CuBiSeCl2 19, 21, 38. Since the MD method includes full-order lattice anharmonicity larger than those lattice anharmonicity considered in BTE up to fourth-order, the results are relatively smaller than those obtained from BTE. Fig. 4(d) reveals that the average $\kappa_{p,SCPH}^{3,4ph}$ at 300 K is 0.584 W m-1 K-1, larger than 0.466 W m-1 K-1 from HNEMD. The latter HNEMD result is a little larger than the experimental $\kappa_{L}$, which is 0.275 W m-1 K-1 at 300 K. This could result from their non-dense sample, with a density of only 88% 3, 39. The result suggests that the higher than fourth-order lattice anharmonicity could be crucial for capturing the ultralow $\kappa_{L}$ of complex compounds such as CuBiSeCl2.
Conclusion
In summary, this study explores the thermal transport properties of CuBiSeCl2 by utilizing a combination of self-consistent phonon theory based on machine learning potentials, first-principles calculations, and a unified approach of the Wigner transport equation to account for both coherent and population contributions to thermal transport. The neuroevolution potentials achieve a level of accuracy comparable to DFT in predicting energies and atomic forces. The temperature-dependent force constants reveal a significant phonon hardening effect over nearly the entire frequency spectrum, indicating strong lattice anharmonicity in CuBiSeCl2. This leads to ultralow $\kappa_{L}$ of 0.81 W m-1 K-1 at 300 K. By using the HNEMD method, where full-order lattice anharmonicity is considered, the $\kappa_{L}$ decreases further to 0.47 W m-1 K-1, closely aligning with experimental measurements 3, 39.
$\kappa_{c}/\kappa_{p}$ exceeds 0.5 above 400 K, indicating that the off-diagonal components of heat flux operators play a significant role in the thermal transport of CuBiSeCl2. Most strikingly, we trace ultralow $\kappa_{L}$ of CuBiSeCl2 to the delocalization of loosely bound copper atoms. Our work may shed light on further suppressing $\kappa_{L}$ and seek for promising thermoelectrics by adjusting bonding strength, localization, and delocalization of electrons 26, 40.
Author Information
Corresponding Author
E-mail: zhibin.gao@xjtu.edu.cn
{acknowledgement}
The authors gratefully acknowledge discussions with Jiongzhi Zheng. We acknowledge the support from the National Natural Science Foundation of China (No.12104356 and No.52250191), and the Fundamental Research Funds for the Central Universities. This work is sponsored by the Key Research and Development Program of the Ministry of Science and Technology (No.2023YFB4604100). We also acknowledge the support by HPC Platform, Xi’an Jiaotong University.

### v3

@@
 Introduction
-Mixed-anion compounds, such as halide-chalcogenides, have recently gained attention as promising candidates for thermoelectric and photovoltaic applications due to their ability to achieve low thermal conductivity while maintaining favorable electronic properties 1, 2.
-These compounds exhibit increased structural anisotropy, dimensional changes, and the formation of novel structural types 3.
-The structural modifications induced by variations in bonding promote the emergence of new physical phenomena and offer opportunities to tune both optoelectronic and thermoelectronic properties by altering the electronic structure 4, 5, 6, 7, 8.
+Mixed-anion compounds, such as halide-chalcogenides, have recently gained attention as promising candidates for thermoelectric and photovoltaic applications due to their ability to achieve low thermal conductivity while maintaining favorable electronic properties 1, 2, 3.
+These compounds exhibit increased structural anisotropy, dimensional changes, and the formation of novel structural types 4.
+The structural modifications induced by variations in bonding promote the emergence of new physical phenomena and offer opportunities to tune both optoelectronic and thermoelectronic properties by altering the electronic structure 5, 6, 7, 8, 9.
 CuBiSeCl2, a mixed-metal chalcohalide, stands out for its potential to exhibit both low thermal conductivity and favorable electronic properties.
-Previous experimental studies have shown that CuBiSeCl2 has a band gap of 1.33 eV, as measured by diffuse reflectance UV-vis spectroscopy, making it suitable for photovoltaic applications 3.
-Additionally, it demonstrates a low thermal conductivity of 0.27 W m-1 K-1 at room temperature 3.
+Previous experimental studies have shown that CuBiSeCl2 has a band gap of 1.33 eV, as measured by diffuse reflectance UV-vis spectroscopy, making it suitable for photovoltaic applications 4.
+Additionally, it demonstrates a low thermal conductivity of 0.27 W m-1 K-1 at room temperature 4.
 However, the complex crystal structure has limited comprehensive studies on the underlying mechanisms responsible for its low $\kappa_{L}$ , particularly for such intricate mixed-anion compounds.
@@
 Our results not only elucidate the origins of the low $\kappa_{L}$ in CuBiSeCl2 but also offer a new strategy for further reducing $\kappa_{L}$ through the delocalization of atoms and the introduction of softer chemical and vibrational bonds.
-The computational details can be found in the Supplemental Material.
+Computational Methods
+A. Structural optimization
+The Vienna Ab initio Simulation Package (VASP) was used to optimize the crystal structure of CuBiSeCl2 10.
+All first-principles calculations are performed by using the PAW method with PBE exchange-correlation functional 11.
+The cut-off energy of 400 eV was used for the wave functions.
+The van der Waals (vdW) interactions are described by using optB86b functional 12, 13.
+The force and energy convergence thresholds of 10-5 eV/Å  and 10-10 eV respectively, were used for both structural relaxation and self-consistent density functional theory (DFT) calculations.
+The cell and the atoms were fully relaxed in the structural optimization process with the Gamma-centered k-point grids of 5×10×3.
+The crystal structure of CuBiSeCl2 belongs to space group $Pnma$ with number of 62 and the optimized lattice constants, using PBE potentials, are $a$ = 8.73 Å, $b$ = 4.05 Å, and $c$ = 12.99 Å, in which the van der Waals (vdW) interactions are described by the optB86b functional 12, 13.
+These values are consistent with the experimental values, which are $a$ = 8.78 Å, $b$ = 4.00 Å, and $c$ = 13.14 Å 4.
+As illustrated in Fig.
+1(b), the phonon spectrum of the PBE functional is free of imaginary frequencies, indicating the dynamical stability of CuBiSeCl2.
+The primitive cell comprises 20 atoms: four Bi, four Se, four Cu, and eight Cl respectively.
+B. Data Generation
+The dataset was generated using random perturbations with the dpdata package 14, 15.
+The deformation ratio of the unit cell was set to 4%, with perturbations in the diagonal part uniformly distributed within the range [-4%, 4%], and perturbations in the off-diagonal part uniformly distributed within the range [-2%, 2%].
+The maximum atomic displacement radius was set to 0.4 Å, meaning that the atomic positions were randomly and uniformly distributed within a spherical space of radius 0.4 Å.
+A total of 600 structures were generated, of which 1/12 were randomly selected as the validation set, while the remaining were used as the training set.
+The structures of CuBiSeCl2 in the dataset were expanded by a 2×4×2 supercell which contains 320 atoms, and the k-point sampling was set to 1×1×1.
+The van der Waals (vdW) interactions is also included.
+C. Training machine learning potential
+To describe the interatomic interaction, machine learning neuroevolution potentials (NEP) was trained on the dataset generated by the DFT calculation 16, 17.
+The NEP parameters can be fitted by minimizing the function 16,
+$L(z)=$ $\lambda_{e}\left(\frac{1}{N_{\text{str}}}\sum_{n=1}^{N_{\text{str}}}\left(U^{\text{NEP}}(n,z)-U^{\text{tar}}(n)\right)^{2}\right)^{1/2}$ (1)
+$+\lambda_{f}\left(\frac{1}{3N}\sum_{i=1}^{N}\left(F_{i}^{\text{NEP}}(z)-F_{i}^{\text{tar}}\right)^{2}\right)^{1/2}$
+$+\lambda_{v}\left(\frac{1}{6N_{\text{str}}}\sum_{n=1}^{N_{\text{str}}}\sum_{\mu\nu}\left(W_{\mu\nu}^{\text{NEP}}(n,z)-W_{\mu\nu}^{\text{tar}}(n)\right)^{2}\right)^{1/2}$
+$+\lambda_{1}\frac{1}{N_{\text{par}}}\sum_{n=1}^{N_{\text{par}}}|z_{n}|$
+$+\lambda_{2}\left(\frac{1}{N_{\text{par}}}\sum_{n=1}^{N_{\text{par}}}z_{n}^{2}\right)^{1/2},$
+where $N_{\text{str}}$ is the number of structures in the training dataset and $N$ is the total number of atoms in all these structures.
+The first three terms are the weighted averages of the energy, forces, and virial tensor using NEP compared with the DFT results.
+That is, the loss functions for energy, force, and virial are defined as their root mean square errors (RMSEs) between the current NEP predictions and the target values.
+The last two terms represent $\ell_{1}$ and $\ell_{2}$ regularization.
+The weights $\lambda_{e}$, $\lambda_{f}$, $\lambda_{v}$, $\lambda_{1}$, and $\lambda_{2}$ are tunable hyperparameters.
+When calculating the loss function, we use the following units: eV/atom for energy and virial, and eV/Å for force components.
+Since the cutoff values for cubic and quartic interactions used in the phonon thermal conductivity calculations are 4.12 Å and 3.30 Å, both smaller than the NEP fitting radial cutoff radius 7.0 Å and angular cutoff radius 7.0 Å, indicating the NEP fitting cutoff radius is sufficiently large for the thermal conductivity calculation.
+Furthermore, we set the number of basis functions that are used to build the radial and angular descriptor to 14, which is greater than the default value of 8, ensuring sufficient accuracy.
+D. Phonon scattering properties
+Three-phonon scattering rates can be expressed as:
+$\Gamma^{+}_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}=$ $\frac{\hbar\pi}{4}\frac{f_{0}^{\prime}-f_{0}^{\prime\prime}}{\omega_{\lambda}\omega_{\lambda^{\prime}}\omega_{\lambda^{\prime\prime}}}\lvert V^{+}_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}\rvert^{2}$ (2)
+$\delta\left(\omega_{\lambda}+\omega_{\lambda^{\prime}}-\omega_{\lambda^{\prime\prime}}\right)$
+$\Gamma^{-}_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}=$ $\frac{\hbar\pi}{4}\frac{f_{0}^{\prime}+f_{0}^{\prime\prime}+1}{\omega_{\lambda}\omega_{\lambda^{\prime}}\omega_{\lambda^{\prime\prime}}}\lvert V^{-}_{\lambda\lambda^{\prime}\lambda^{\prime\prime}}\rvert^{2}$ (3)
+$\delta\left(\omega_{\lambda}-\omega_{\lambda^{\prime}}-\omega_{\lambda^{\prime\prime}}\right)$
+where, for simplicity, $f_{0}^{\prime}$ stands for $f_{0}\bigl{(}\omega_{\lambda^{\prime}}\bigr{)}$ and $f_{0}^{\prime\prime}$ denotes $f_{0}\bigl{(}\omega_{\lambda^{\prime\prime}}\bigr{)}$.
+The term $\Gamma_{\lambda\lambda^{\prime},\lambda^{\prime\prime}}^{+}$ corresponds to an absorption process, where two incident phonons $(\omega_{\lambda},\omega_{\lambda^{\prime}})$ merge into a single phonon whose frequency is their sum.
+Meanwhile, $\Gamma_{\lambda\lambda^{\prime},\lambda^{\prime\prime}}^{-}$ describes an emission process, in which one phonon splits into two $(\omega_{\lambda^{\prime}},\omega_{\lambda^{\prime\prime}})$ 18.
+Phonon density of states (DOS) is calculated with a linear tetrahedron method which is default in Phonopy 19.
+Projected DOS is calculated using “PDOS = AUTO”.
+Without specifying projection direction, PDOS is computed as sum of $g^{j}(\omega,\hat{\mathbf{n}})$ projected onto cartesian axes $x,y,z$,
+$g^{j}(\omega)=\sum_{\hat{\mathbf{n}}=\{x,y,z\}}g^{j}(\omega,\hat{\mathbf{n}}).$ (4)
+where $j$ is the atom indices and $\hat{\mathbf{n}}$ is the unit projection direction vector.
+Mean square displacement of different atoms is also calculated using Phonopy package 19. The expectation value of the squared atomic displacement is calculated as,
+$\left\langle|u^{\alpha}(jl,t)|^{2}\right\rangle=$ $\frac{\hbar}{2Nm_{j}}\sum_{\mathbf{q},\nu}\omega_{\nu}(\mathbf{q})^{-1}$ (5)
+$\left(1+2n_{\nu}(\mathbf{q},T)\right)|e_{\nu}^{\alpha}(j,\mathbf{q})|^{2},$
+where $n_{\nu}(\mathbf{q},T)$ is the phonon population, which is given by,
+$n_{\nu}(\mathbf{q},T)=\frac{1}{\exp(\hbar\omega_{\nu}(\mathbf{q})/k_{\text{B}}T)-1},$ (6)
+$T$ is the temperature, and $k_{\text{B}}$ is the Boltzmann constant.
+Atomic participation ratio (APR) is calculated with alamode package and is defined as follow:
+$APR_{\mathbf{q}j,\kappa}$ $=\frac{|\mathbf{e}(\kappa;\mathbf{q}j)|^{2}}{M_{\kappa}}$ (7)
+$\quad/\left(N_{\kappa}\sum_{\kappa}^{N_{\kappa}}\frac{|\mathbf{e}(\kappa;\mathbf{q}j)|^{4}}{M_{\kappa}^{2}}\right)^{1/2}$
+APR is an atomic decomposition of PR that satisfies $PR_{\mathbf{q}j}=\sum_{\kappa}(APR_{\mathbf{q}j,\kappa})^{2}$.
+E. Lattice thermal conductivity
+The lattice thermal conductivity $\kappa_{p}$ based on Peierls-Boltzmann transport theory is given by20,
+$\kappa_{p}=\frac{\hbar^{2}}{k_{B}T^{2}VN_{0}}\sum_{\lambda}n_{\lambda}(n_{\lambda}+1)\omega^{2}_{\lambda}\bm{v}\lambda\otimes\bm{v}\lambda\bm{\tau}_{\lambda},$ (8)
+where $\hbar$ denotes the reduced Planck constant, $k_{B}$ the Boltzmann constant, $T$ the absolute temperature, $V$ the volume of the primitive unit cell, and $N_{0}$ the number of phonon wave vectors sampled in the first Brillouin zone.
+The variables $n_{\lambda}$, $\omega_{\lambda}$, $\bm{v}_{\lambda}$, and $\tau_{\lambda}$ indicate the phonon population, frequency, group velocity, and lifetime for the $\lambda$ mode which is defined by wave vector $q$ and branch index $s$.
+In the interatomic force constants (IFCs) calculations, a 2$\times$4$\times$1 supercell of CuBiSeCl2 was used.
+Third- and fourth-order IFCs were calculated up to the eighth and third nearest neighbors, respectively 21, 22.
+The q-mesh grid was set to 6$\times$12$\times$4 for four-phonon (4ph) processes and 10$\times$20$\times$7 for three-phonon (3ph) processes.
+The self-consistent phonon (SCPH) approximation 23, 24 was applied to incorporate first-order phonon frequency corrections due to quartic anharmonicity, which enhances accuracy for materials with soft phonon modes and significant anharmonicity.
+The SCPH is described by
+${\Omega}_{\lambda}^{2}={\omega}_{\lambda}^{2}+2{\Omega}_{\lambda}\sum\limits_{{\lambda}_{1}}I_{\lambda\lambda_{1}},$ (9)
+where $\omega_{\lambda}$ represents the initial phonon frequency derived from the harmonic approximation, and $\Omega_{\lambda}$ denotes the temperature-dependent renormalized phonon frequency.
+The scalar $I_{\lambda\lambda_{1}}$ is calculated as follows:
+${I_{\lambda\lambda_{1}}}=\frac{\hbar}{8N_{0}}\frac{V^{(4)}(\lambda,-\lambda,\lambda_{1},-\lambda_{1})}{\Omega_{\lambda}\Omega_{\lambda_{1}}}\left[1+2n_{\lambda}(\Omega_{\lambda_{1}})\right],$ (10)
+where $V^{(4)}$ represents the fourth-order IFCs in the reciprocal space, and the phonon population $n_{\lambda}$ follows the Bose-Einstein distribution.
+Eq.
+(9) and Eq.
+(10) both possess the parameters $I_{\lambda\lambda_{1}}$ and $\Omega_{\lambda}$, allowing the SCPH equation to be solved iteratively.
+Also, we fix the volume in self-consistent phonon calculation which means that thermal expansion was not considered in temperature-dependent renormalized phonon frequency 25.
+To enhance the accuracy of the lattice thermal conductivity, the coherent phonon $\kappa_{c}$ describing tunneling of coherent phonons, was also included and is defined as26
+$\kappa_{c}=$ $\frac{\hbar^{2}}{k_{B}T^{2}VN_{0}}\sum_{\bm{q}}\sum_{s\neq s^{\prime}}\frac{\omega_{\bm{q}}^{s}+\omega_{\bm{q}}^{s^{\prime}}}{2}\bm{v}_{\bm{q}}^{s,\bm{s}^{\prime}}\bm{v}_{\bm{q}}^{\bm{s}^{\prime},s}$ (11)
+×$\times\frac{\omega_{\bm{q}}^{s}n_{\bm{q}}^{s}\left(n_{\bm{q}}^{s}+1\right)+\omega_{\bm{q}}^{s^{\prime}}n_{\bm{q}}^{s^{\prime}}\left(n_{\bm{q}}^{s^{\prime}}+1\right)}{4\left(\omega_{\bm{q}}^{s^{\prime}}-\omega_{\bm{q}}^{s}\right)^{2}+\left(\Gamma_{\bm{q}}^{s}+\Gamma_{\bm{q}}^{s^{\prime}}\right)^{2}}$
+×$\times\left(\Gamma_{\bm{q}}^{s}+\Gamma_{\bm{q}}^{s^{\prime}}\right),$
+where $\hbar$ is the reduced Planck constant, $k_{B}$ is Boltzmann constant, $V$ is the primitive cell volume, and $N_{0}$ is the total number of sampled phonon vectors.
+Additionally, the temperature-dependent $\kappa_{L}$ of CuBiSeCl2 is also calculated from homogeneous nonequilibrium molecular dynamics (HNEMD) simulations 27, 28, 29, an average result of 2ns for 5 times with the model cell around 75 Å along each edge.
+The total heat flow in the system is the sum of contributions from each atom,
+$J=\sum_{i}\sum_{j\neq i}\mathbf{r}_{ij}\left(\frac{\partial U_{j}}{\partial\mathbf{r}_{ji}}\cdot\mathbf{v}_{i}\right)$ (12)
+where J in the Green–Kubo method for calculating thermal conductivity, with ${r}_{ij}$ is the distance vector between atoms i and j, U is the potential energy, and v is velocity.
+In the HNEMD method, each atom in the system is subjected to an external force to drive it out of equilibrium.
+This force, expressed as
+$\vec{F}_{\text{ext},i}=E_{i}\vec{F}_{e}+\vec{F}_{e}\cdot\vec{W}_{i},$ (13)
+This approach offers a way to compute thermal conductivity.
+Unlike the NEMD method, the HNEMD framework maintains the system in a uniform non-equilibrium state, without explicit heat sources or sinks.
+As a result, phonon boundary scattering is absent, and heat circulates due to periodic boundary conditions along the transport direction.
+This method produces finite-size effects similar to those seen in the EMD approach.
 Results and discussion
-A. Crystal structure
 Figure 1: (a) The conventional cell structure of CuBiSeCl2.
@@
 (f) Calculated spectral and cumulative particle-like lattice thermal conductivity ($\kappa_{p}$) of CuBiSeCl2 using the SCPH+3,4ph model at 300 K in different axes.
-The crystal structure of CuBiSeCl2 belongs to space group $Pnma$ with number of 62 and the optimized lattice constants, using PBE potentials, are $a$ = 8.73 Å, $b$ = 4.05 Å, and $c$ = 12.99 Å, in which the van der Waals (vdW) interactions are described by the optB86b functional 9, 10.
-These values are consistent with the experimental values, which are $a$ = 8.78 Å, $b$ = 4.00 Å, and $c$ = 13.14 Å 3.
-As illustrated in Fig.
-1(b), the phonon spectrum of the PBE functional is free of imaginary frequencies, indicating the dynamical stability of CuBiSeCl2.
-The primitive cell comprises 20 atoms: four Bi, four Se, four Cu, and eight Cl respectively.
-B. Anharmonic lattice dynamics
+A. Anharmonic lattice dynamics
 The impact of fourth-order anharmonicity on the phonon spectra at finite temperatures is depicted in Fig.
@@
 The overall phonon hardening is observed in crystalline CuBiSeCl2 as the temperature increases from 0 to 700 K.
-This suggests the presence of strong lattice anharmonicity in CuBiSeCl2 and highlights the importance of properly treating temperature renormalization 11, 12, 13, 14, 15.
+This suggests the presence of strong lattice anharmonicity in CuBiSeCl2 and highlights the importance of properly treating temperature renormalization 30, 23, 31, 32, 33.
 In particular, the low-frequency phonons below 3 Thz are predominantly influenced by the Cu (yellow) and Bi (red) atoms, as evidenced by the projected phonon density of states (PDOS) in Fig.
@@
 1(d-e) show that Cu and Bi atoms contribute mostly to the low-lying flattened phonon bands in the frequency lower than 2.0 THz.
-This suggests that strong phonon scattering is occurring, which suppresses the primary heat carriers, i.e., acoustic phonon modes 16, 17, 18.
 In contrast, the high-frequency phonon modes around 6 THz in CuBiSeCl2 are predominantly contributed by Se and Cl atoms and exhibit a temperature-dependent softening in Fig.
@@
 1(f) also approve that low frequency phonons less than 1.5 THz are the main contributor to the $\kappa_{p}$ of CuBiSeCl2.
-C. Delocalization of copper atoms
+These indicate that the phonon vibrations of Cu and Bi atoms predominantly govern $\kappa_{p}$ in CuBiSeCl234, 35, 36.
+B. Delocalization of copper atoms
 Fig.
-2(a) shows the mean square displacements (MSDs) of atoms in CuBiSeCl2 as a function of temperature along the $b$-axis.
+2(a) shows the mean square displacements (MSDs) of atoms obtained by both harmonic approximation (HA) and self-consistent phonon (SCPH) in CuBiSeCl2 as a function of temperature calculations along the $b$-axis.
 MSD is a measure of the deviation of the position of an atom with respect to a reference position over time.
-It indicates the atomic displacements of Cu are significantly larger than those of other atoms Bi, Se, and Cl, suggesting that the bonding of Cu atoms is heavily delocalized and soft 11.
-In Fig.
-2(b), the displacement of Cu atoms shows a relatively deep energy well with a flat bottom as a function of the phonon mode along the $b$-axis.
-The U-shaped potential energy curve significantly deviates from the harmonic approximation (quadratic fit) shown as the dotted line in Fig.
-2(b).
-The U-shaped potential energy of Cu atom displacement can be well fitted by a fourth-order polynomial (solid yellow line in Fig.
-2(b)), expressed as $y=1.41x^{4}+0.38x^{2}$, indicating strong quartic phonon anharmonicity.
+It indicates the atomic displacements of Cu are significantly larger than those of other atoms Bi, Se, and Cl, suggesting that the bonding of Cu atoms is heavily delocalized and soft 30.
+Meanwhile, the MSDs calculated by HA typically exhibits overestimation, while anharmonic effects suppress the vibrational amplitudes of all atoms 30.
+The discrepancy between them can intuitively quantify the role of anharmonic effects.
+Fig.
+2(b) illustrates the differences in MSDs between the HA and SCPH calculations along all three axes for CuBiSeCl2.
+It can be observed that the MSDs differences of Cu atoms are larger than those of other elements in all three directions.
+Particularly, the MSDs difference along $b$-axis is the largest, indicating strong phonon anharmonicity in this direction.
+Besides, the U-shaped potential energy of Cu atom displacement can be well fitted by a fourth-order polynomial in Fig.
+S8 37, also qualitatively indicating strong high-order phonon anharmonicity 38.
 The red color corresponds to complete localization and blue corresponds to complete delocalization of electrons in 2D electron localization function (ELF) of CuBiSeCl2 shown in Fig.
 2(c).
-It can be seen that the localized electronic density around Cu is extremely low, significantly lower than the other atoms, which reveals that the Se and Cl elements exhibit a much stronger attraction of electrons than the Cu elements, suggesting a greater degree of delocalization for Cu atoms.
+The deeper blue color between Cu and Cl or Se atoms indicates there is no significant electron wave function overlap between Cu and its surrounding atoms, further suggesting that no strong bonding is formed and a greater degree of delocalization for Cu atoms.
 The $\kappa_{L}$ of CuBiSeCl2 at 300 K in Fig.
-2(d-f), calculated using homogeneous nonequilibrium molecular dynamics (HNEMD) is the averaged result from five 2 ns simulation runs, totally 10 ns 19, 20, 21.
+2(d-f), calculated using homogeneous nonequilibrium molecular dynamics (HNEMD) is the averaged result from five 2 ns simulation runs, totally 10 ns 27, 28, 29.
 We employed precise NEP to obtain the intrinsic $\kappa_{L}$ of CuBiSeCl2, highlighted in red color in Fig.2(d-f).
@@
 These values indicate a significantly lower intrinsic $\kappa_{L}$ along the $c$-axis (0.34 W m-1 K-1) compared to the $a$- and $b$-axes (0.50 and 0.55 W m-1 K-1), consistent with the final result shown in Fig.
-4(d) 13, 22.
+4(d) 31, 39.
 Based on Fig.
 2(a-c), we observe a significant difference in the delocalization of Cu compared to Bi, Se, and Cl.
-To investigate the impact of the delocalization of different atoms on $\kappa_{L}$, we used the same HNEMD method to calculate $\kappa_{L}$ along $a$-, $b$- and $c$-axes of the system after fixing different types of atoms.
+To investigate the impact of the delocalization of different atoms on $\kappa_{L}$, we used the same HNEMD method to calculate $\kappa_{L}$ along $a$-, $b$- and $c$-axes of the system after fixing different types of atoms by setting the velocities and forces acting on the selected atoms to zero with the function implanted in GPUMD software package 16.
 The results are shown in Fig.
@@
 However, fixing Cu in the $a$- and $b$-axes reveals its counterintuitive phenomenon, as its fixation leads to an increase in $\kappa_{L}$ compared to the intrinsic value by 0.08 and 0.67 W m-1 K-1, respectively.
-This highlights that the delocalization of Cu has a negative impact on $\kappa_{L}$, especially on the $b$-axis 23.
+This highlights that the delocalization of Cu has a negative impact on $\kappa_{L}$, especially on the $b$-axis 40.
 Figure 2: (a) The mean square atomic displacements (MSDs) with temperatures for CuBiSeCl2.
-(b) Potential energy per atom with displacement of Cu (yellow) atoms along the $b$-axis direction with quartic and quadratic polynomial fitting.
-It writes as $y=1.41x^{4}+0.38x^{2}$ and $y=1.58x^{2}$.
-For the Bi (red), Se (blue), and Cl (purple) atoms, the quartic polynomial fittings along the $b$-axis direction are as follows: $y=0.67x^{4}+1.75x^{2}$, $y=1.20x^{4}+1.39x^{2}$ and $y=0.015x^{4}+2.76x^{2}$.
-The cubic term is minimal and less than $10^{-5}$ 24, 22.
+The solid and dashed lines correspond to the MSDs obtained by harmonic approximation (HA) and self-consistent phonon (SCPH) calculations respectively.
+(b) Difference between MSDs obtained by HA and SCPH approaches in $a$-axis, $b$-axis and $c$-axis.
 (c) Electron localization function (ELF) of CuBiSeCl2.
@@
 The values labeled on the right represent the averaged final converged $\kappa_{L}$.
-D. Thermal transport properties
-Based on Peierls-Boltzmann transport theory, the lattice thermal conductivity $\kappa_{p}$ is given by,
-$\kappa_{p}=\frac{\hbar^{2}}{k_{B}T^{2}VN_{0}}\sum_{\lambda}n_{\lambda}(n_{\lambda}+1)\omega^{2}_{\lambda}\bm{v}\lambda\otimes\bm{v}\lambda\bm{\tau}_{\lambda},$ (1)
-and $\kappa_{c}$ describing tunneling and wavelike of coherent phonons, governed by the Wigner transport equation, is defined as 25, 13,
-$\kappa_{c}=$ $\frac{\hbar^{2}}{k_{B}T^{2}VN_{0}}\sum_{\bm{q}}\sum_{s\neq s^{\prime}}\frac{\omega_{\bm{q}}^{s}+\omega_{\bm{q}}^{s^{\prime}}}{2}\bm{v}_{\bm{q}}^{s,\bm{s}^{\prime}}\bm{v}_{\bm{q}}^{\bm{s}^{\prime},s}$ (2)
-×$\times\frac{\omega_{\bm{q}}^{s}n_{\bm{q}}^{s}\left(n_{\bm{q}}^{s}+1\right)+\omega_{\bm{q}}^{s^{\prime}}n_{\bm{q}}^{s^{\prime}}\left(n_{\bm{q}}^{s^{\prime}}+1\right)}{4\left(\omega_{\bm{q}}^{s^{\prime}}-\omega_{\bm{q}}^{s}\right)^{2}+\left(\Gamma_{\bm{q}}^{s}+\Gamma_{\bm{q}}^{s^{\prime}}\right)^{2}}$
-×$\times\left(\Gamma_{\bm{q}}^{s}+\Gamma_{\bm{q}}^{s^{\prime}}\right),$
-The total $\kappa_{L}$ = $\kappa_{p}$ + $\kappa_{c}$, and the parameters in these equations can be found in the Supplementary Material.
+C. Thermal transport properties
 From Eq.
@@
 3(a), with temperature going higher than 300 K, the $C_{V}$ given out by DFT calculation is very close to the theoretical value $3Nk_{B}$ = 0.295 J g-1 K-1.
-Also, this value of CuBiSeCl2 is even lower than 0.36 J g-1 K-1 of liquid-like thermoelectric Cu2Se 26 and a little higher than 0.156 J g-1 K-1 of renowned PbTe 27.
+Also, this value of CuBiSeCl2 is even lower than 0.36 J g-1 K-1 of liquid-like thermoelectric Cu2Se 41 and a little higher than 0.156 J g-1 K-1 of renowned PbTe 42.
 […]
@@
 This is closely linked to the Cu and Bi atoms and the bonding hierarchy.
-Moreover, the total Grüneisen value is 1.50, close to PbTe value of 1.45 31, indicating CuBiSeCl2 exhibits substantial anharmonicity.
+Moreover, the total Grüneisen value is 1.50, close to PbTe value of 1.45 46, indicating CuBiSeCl2 exhibits substantial anharmonicity.
 To deepen the analysis of phonon scattering processes, we computed the phonon scattering phase space, which evaluates the channel available for phonon-phonon scattering.
@@
 3(e).
-The scattering process of the AAO mode is found to be as strong as the AOO mode below 1 THz, meaning that the AAO and AOO phonons play dominant roles in heat transport.
+The scattering process of the AAO mode is found to be as strong as the AOO mode below 1 THz, meaning that the AAO and AOO phonons play important roles in reduce heat transport.
 Also, as temperature increases to 700 K in Fig.
@@
 (f) Phonon lifetime of at 300 K.
-The red horizontal line represents the Wigner limit, expressed as $\tau_{\text{threshold}}=[\Delta\omega_{\text{avg}}]^{-1}$, where $\Delta\omega_{\text{avg}}$ is the average phonon inter-band spacing 32.
+The red horizontal line represents the Wigner limit, expressed as $\tau_{\text{threshold}}=[\Delta\omega_{\text{avg}}]^{-1}$, where $\Delta\omega_{\text{avg}}$ is the average phonon inter-band spacing 47.
 $\Delta\omega_{\text{avg}}$ is determined by $\omega_{\text{max}}/3N_{\text{at}}$, with $N_{\text{at}}$ being the number of atoms per unit cell and $\omega_{\text{max}}$ being the highest frequency.
-Phonons with lifetimes longer than $\tau_{\text{threshold}}$ primarily affect thermal conductivity via population mechanisms, while shorter lifetimes contribute to wavelike tunneling.
+Phonons with lifetimes longer than $\tau_{\text{threshold}}$ primarily affect thermal conductivity via population mechanisms (green region), while shorter lifetimes contribute to wavelike tunneling (blue region).
 The black curve shows the Ioffe-Regel limit, $\tau_{\text{IR}}=2\pi/\omega$, in which $\omega$ is the phonon frequency.
-Phonons lifetime exceeding $\tau_{\text{IR}}$ can be described by the Wigner transport equation 33, 15, 22.
-Recently, a seminal concept is introduced known as the Wigner limit in time, which has been applied to distinguish phonons into various thermal transport regimes, primarily consisting of particle-like propagation and wavelike tunneling 32, 25, 34.
+Phonons lifetime exceeding $\tau_{\text{IR}}$ can be described by the Wigner transport equation 48, 33, 39.
+Recently, a seminal concept is introduced known as the Wigner limit in time, which has been applied to distinguish phonons into various thermal transport regimes, primarily consisting of particle-like propagation and wavelike tunneling 47, 26, 49.
 The Wigner limit in time is characterized as $\tau_{\text{threshold}}=3N_{\text{at}}/\omega_{\text{max}}$, where $N_{\text{at}}$ represents the number of atoms in the unit cell, and $\omega_{\text{max}}$ denotes the maximum phonon frequency.
@@
 4(b).
-These values represent reductions of 17%, 33%, and 42% compared to $\kappa_{p,SCPH}^{3ph}$, highlighting the impact of 4ph scattering 13, 18, 35.
+These values represent reductions of 17%, 33%, and 42% compared to $\kappa_{p,SCPH}^{3ph}$, highlighting the impact of 4ph scattering 31, 36, 50.
 This result confirms that phonon hardening and 4ph scattering, driven by quartic anharmonicity, have counteractive effects on $\kappa_{L}$.
@@
 (c) $\kappa_{p}$ of CuBiSeCl2 compare with $\kappa_{c}$.
-(d) $\kappa_{L}$ of CuBiSeCl2 using BTE method and MD method compare with experimental values 3.
+(d) $\kappa_{L}$ of CuBiSeCl2 using BTE method and MD method compare with experimental values 4.
 Nevertheless, the coherence contributions are not negligible as shown in Fig.
 4(c).
-By further considering the off-diagonal terms of heat flux operators, the total $\kappa_{L}$ significantly enhanced by incorporating both $\kappa_{p,SCPH}^{3,4ph}$ and $\kappa_{c}$ 25.
+By further considering the off-diagonal terms of heat flux operators, the total $\kappa_{L}$ significantly enhanced by incorporating both $\kappa_{p,SCPH}^{3,4ph}$ and $\kappa_{c}$ 26.
 The ratio $\kappa_{c}/\kappa_{p}$ (red star) is greater than 0.5 at 500 K and exhibits a pronounced increase with temperature.
@@
 This observation suggests that the contribution of $\kappa_{c}$ is equally important as $\kappa_{p}$ in crystalline CuBiSeCl2 especially in $a$-axis and high temperature.
-The dominant role of coherence contribution in thermal transport was also observed in other complex compounds such as Cs2AgBiBr6 33, Cu12Sb4S13 36, CsPbBr3 37, TlAgI2 15, La2Zr2O7 32, Tl9SbTe6 22, and even simple cubic CsCl 35.
-We further use the HNEMD method to assess the intrinsic $\kappa_{L}$ of CuBiSeCl2 19, 21, 38.
+The dominant role of coherence contribution in thermal transport was also observed in other complex compounds such as Cs2AgBiBr6 48, Cu12Sb4S13 51, CsPbBr3 52, TlAgI2 33, La2Zr2O7 47, Tl9SbTe6 39, and even simple cubic CsCl 50.
+We further use the HNEMD method to assess the intrinsic $\kappa_{L}$ of CuBiSeCl2 27, 29, 16.
 Since the MD method includes full-order lattice anharmonicity larger than those lattice anharmonicity considered in BTE up to fourth-order, the results are relatively smaller than those obtained from BTE.
@@
 The latter HNEMD result is a little larger than the experimental $\kappa_{L}$, which is 0.275 W m-1 K-1 at 300 K.
-This could result from their non-dense sample, with a density of only 88% 3, 39.
+This could result from their non-dense sample, with a density of only 88% 4, 53, 54, 37.
 The result suggests that the higher than fourth-order lattice anharmonicity could be crucial for capturing the ultralow $\kappa_{L}$ of complex compounds such as CuBiSeCl2.
 Conclusion
-In summary, this study explores the thermal transport properties of CuBiSeCl2 by utilizing a combination of self-consistent phonon theory based on machine learning potentials, first-principles calculations, and a unified approach of the Wigner transport equation to account for both coherent and population contributions to thermal transport.
+In summary, this study explores the thermal transport properties of CuBiSeCl2 by utilizing a combination of self-consistent phonon theory based on machine learning potentials 55, 16, 17, first-principles calculations, and a unified approach of the Wigner transport equation to account for both coherent and population contributions to thermal transport.
 The neuroevolution potentials achieve a level of accuracy comparable to DFT in predicting energies and atomic forces.
@@
 This leads to ultralow $\kappa_{L}$ of 0.81 W m-1 K-1 at 300 K.
-By using the HNEMD method, where full-order lattice anharmonicity is considered, the $\kappa_{L}$ decreases further to 0.47 W m-1 K-1, closely aligning with experimental measurements 3, 39.
+By using the HNEMD method, where full-order lattice anharmonicity is considered, the $\kappa_{L}$ decreases further to 0.47 W m-1 K-1, closely aligning with experimental measurements 4, 53.
 $\kappa_{c}/\kappa_{p}$ exceeds 0.5 above 400 K, indicating that the off-diagonal components of heat flux operators play a significant role in the thermal transport of CuBiSeCl2.
 Most strikingly, we trace ultralow $\kappa_{L}$ of CuBiSeCl2 to the delocalization of loosely bound copper atoms.
-Our work may shed light on further suppressing $\kappa_{L}$ and seek for promising thermoelectrics by adjusting bonding strength, localization, and delocalization of electrons 26, 40.
+Our work may shed light on further suppressing $\kappa_{L}$ and seek for promising thermoelectrics by adjusting bonding strength, localization, and delocalization of electrons 41, 56, 57..
 Author Information
@@
 The authors gratefully acknowledge discussions with Jiongzhi Zheng.
-We acknowledge the support from the National Natural Science Foundation of China (No.12104356 and No.52250191), and the Fundamental Research Funds for the Central Universities.
+We acknowledge the support from the National Natural Science Foundation of China (No.52250191), and the Fundamental Research Funds for the Central Universities.
 This work is sponsored by the Key Research and Development Program of the Ministry of Science and Technology (No.2023YFB4604100).
 We also acknowledge the support by HPC Platform, Xi’an Jiaotong University.
+Data Availability
+The structure data of CuBiSeCl2 that we used can be found in Supplemental Material 37. Other data is available from the authors upon reasonable request.